OpenStudy (anonymous):

Convert the integral from rectangular coordinates to cylindrical coordinates.

OpenStudy (anonymous):

\[\int\limits_{-2}^{2}\int\limits_{-\sqrt{4-x^2}}^{\sqrt{4-x^2}}\int\limits_{x^2+y^2}^{4}xdzdydx\]